A rivet is to be inserted into a hole. If the Standard deviation of hole diameter exceeds


0,01mm, there is an unacceptably high probability that the rivet will not fit. A random sample


of n=15 part is selected and the hole diameter is measured. The sample Standard deviation of


the hole diameter measurements is s=0.008mm. Is there strong evidence to indicate that the


standard deviation of hole diameter exceeds 0.01mm? Use α=0.01

1
Expert's answer
2022-01-24T14:49:56-0500
"\sigma_0=0.01\ mm, \sigma_0^2=0.0001\ mm^2"

"s=0.008\ mm, s^2=0.000064\ mm^2"

The following null and alternative hypotheses need to be tested:

"H_0:\sigma^2\leq0.0001"

"H_1:\sigma^2>0.0001"

This corresponds to a right-tailed test test, for which a Chi-Square test for one population variance will be used.

Based on the information provided, the significance level is "\alpha = 0.01," "df=n-1=15-1=14" degrees of freedom, and the the rejection region for this right-tailed test is "R = \{\chi^2: \chi^2 >29.1412\}."

The Chi-Squared statistic is computed as follows:


"\chi^2=\dfrac{(n-1)s^2}{\sigma_0^2}=\dfrac{(15-1)(0.000064)}{0.0001}=8.96"

Since it is observed that "\chi^2 = 8.96 \le 29.1412= \chi_c^2," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population variance "\sigma^2" is greater than "0.0001," at the "0.01" significance level.

Therefore, there is not enough evidence to claim that the standard deviation of hole diameter exceeds "0.01\ mm" at the "0.01" significance level.